A man invested $1000, part at 5% and the rest at 6%. His total annual income was $53. How much did he invest at each rate?
A man invested $2500, part at 4% and part at 5%. His total annual income was $110. How much did he invest at each rate?
A girl has ten coins, all nickels and dimes, with a total value of 70 cents. Find the number of each kind of coin.
A man spends $1.60 for thirty-four stamps, getting some 5 cent stamps and some 4 cent stamps . How much of each kind does he buy?

Answer by checkley77(12569)

.06x+.05(1000-x)=53
.06x+50-.05x=53
.01x=53-50
.01x=3
x=3/.01
x=300 dollars invested @ 6%
1000-300=700 dollars invested @ 5%
proof
.06*300+.05*700=53
18+35=53
53=53